### v2

# Gravitational Positivity Bounds on Higgs-Portal Light Dark Matter

Abstract
Gravitational positivity bounds are constraints on a renormalizable theory in the presence of a massless graviton, under the assumption that the gravitational theory is ultraviolet-completed by a perturbative string theory. We derive these bounds for the Higgs-portal scalar dark matter model using the forward scattering process $\phi\phi\to\phi\phi$. We find that, in the absence of a dark matter self-coupling, new physics beyond the Higgs-portal dark matter interaction must appear below an energy scale of $10^{10}$ GeV if the dark matter mass is smaller than the Higgs boson mass. The presence of a dark matter self-coupling alters this situation. A hierarchy between the dark matter four-point self-coupling $\lambda_{\phi}$ and a tiny Higgs-portal coupling $\lambda_{h\phi}$ is required to raise the energy scale at which the new physics appears. If $\lambda_{\phi}/\lambda_{h\phi}=10^{12}$, the dark matter model can remain valid up to the grand unified theory (GUT) scale or the typical string scale. In this case, the relic abundance of dark matter in the Universe can be reproduced via the dark freeze-out scenario. A parameter set with $\lambda_{\phi}\sim\mathcal{O}(1)$, $\lambda_{h\phi}\sim 10^{-12}$, and a sub-GeV dark matter mass can accommodate the GUT-scale $\Lambda$ within the Higgs-portal light dark matter framework.
1 Introduction
The existence of dark matter (DM) is one of the primary motivations for considering physics beyond the Standard Model (SM). If we assume the existence of a SM gauge-singlet scalar particle, it can interact with the Higgs boson through a renormalizable Lagrangian. When this scalar field is odd under a $Z_{2}$ symmetry, whereas all SM particles are even, the scalar is stable and does not decay into SM particles. As a result, the scalar boson becomes a viable DM candidate, and its interaction with all other particles proceeds through the so-called Higgs portal. This Higgs-portal scalar DM model is a minimal and well-studied extension of the SM with rich phenomenology [55, 44, 20] (see also [15, 41] for reviews and [38, 39, 58, 26] for studies including higher-dimensional operators).
The viability of such DM models can be evaluated through their ability to reproduce the observed DM relic abundance in the Universe, and through their consistency with experimental and observational constraints. In addition to these phenomenological approaches, the models can be examined from the perspective of ultraviolet (UV) completion, taking into account the fundamental principles of unitarity, analyticity, and Lorentz invariance of the $S$-matrix.
A powerful tool in this regard is the use of positivity bounds [3, 51, 8]. The original positivity bounds were derived for higher-dimensional operators of dimension-8 and above. Dimension-8 operators are accompanied by the fourth power of the cutoff scale in the denominator, so they are suppressed compared to dimension-6 and dimension-5 operators, or the unsuppressed renormalizable interactions. However, by introducing observables that are insensitive to dimension-6 operators, one can isolate the effects of dimension-8 interactions and explore the imposed positivity bounds and their possible violations [42]. Positivity bounds on Higgs-portal scalar DM with local dimension-8 operators have been investigated by Kim, Lee, and the author in [38, 39].
Constraints can also be placed on renormalizable interactions by exploiting gravitational positivity bounds [56]. Assuming that gravity is UV-completed by a weakly coupled string theory, the renormalizable theory must satisfy bounds imposed by this UV gravitational completion. These constraints are referred to as gravitational positivity bounds [56] (see also [35] for related arguments involving Regge states coupled to the graviton, and [12, 48, 47, 13, 40] for applications to the SM and phenomenological models).
In this paper, we derive gravitational positivity bounds for the renormalizable Higgs-portal scalar DM interactions and discuss their phenomenological implications. This paper is organized as follows. In Section 2, we introduce the Lagrangian of the model. In Section 3, we review the derivation of gravitational positivity bounds for renormalizable interactions and present the bounds obtained for the Higgs-portal scalar DM model. In Section 4, we discuss the phenomenological implications of these bounds for the Higgs-portal real scalar DM model. Finally, in Section 5, we summarize our findings and outline directions for future work.
2 Higgs-Portal Scalar Dark Matter
We consider DM to be a real scalar field $\phi$, which is a SM gauge singlet and odd under a discrete $Z_{2}$ (dark) parity, i.e., $\phi\to-\phi$, whereas all SM particles are even under this parity. We assume that the Lagrangian is symmetric under the $Z_{2}$ transformation to ensure DM stability, so that a DM particle $\phi$ cannot decay into SM particles. The renormalizable interaction of DM with the SM sector is of the Higgs-portal type, meaning that the interactions occur through the Higgs boson. The Lagrangian for the Higgs-portal scalar DM field $\phi$ is given by [37, 18]
$\mathcal{L}=\mathcal{L}_{\mathrm{SM}}+\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}\mu^{2}_{\phi}\phi^{2}-\frac{1}{4!}\lambda_{\phi}\phi^{4}-\frac{1}{2}\lambda_{h\phi}H^{\dagger}H\phi^{2},$ (1)
where $\mathcal{L}_{\mathrm{SM}}$ is the SM Lagrangian and $H$ is the SM Higgs doublet. The bare mass term of the DM field is allowed, whereas linear and cubic terms such as $\phi$, $\phi^{3}$, and $H^{\dagger}H\phi$ are forbidden by the $Z_{2}$ symmetry.
The SM Higgs doublet acquires a vacuum expectation value (VEV) $v$, leading to spontaneous electroweak symmetry breaking. In the unitary gauge, the Higgs doublet is written as
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v+h\end{pmatrix}$, (2)
where $h$ denotes the physical SM Higgs field. Substituting Eq. (2) into Eq. (1), we obtain
$\mathcal{L}=\mathcal{L}_{\mathrm{SM}}+\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{2}m^{2}_{\phi}\phi^{2}-\frac{1}{4!}\lambda_{\phi}\phi^{4}-\frac{1}{2}\lambda_{h\phi}vh\phi^{2}-\frac{1}{4}\lambda_{h\phi}h^{2}\phi^{2},$ (3)
with the physical DM mass given by
$m^{2}_{\phi}=\mu^{2}_{\phi}+\frac{1}{2}\lambda_{h\phi}v^{2}.$ (4)
This model is characterized by three independent parameters: the scalar DM mass (expressed in terms of the Lagrangian parameters via Eq. (4)), the Higgs-portal coupling, and the DM self-coupling, which we denote by
$\{m_{\phi},\lambda_{h\phi},\lambda_{\phi}\}.$ (5)
3 Gravitational Positivity Bounds and Application for Higgs-Portal Scalar DM
Gravitational positivity bounds [56] can be applied to renormalizable interactions. In this section, we first review the framework of gravitational positivity bounds. Following this review, in Sec. 3.2, we derive the gravitational positivity bound for the $\phi\phi\to\phi\phi$ process using the Higgs-portal scalar DM Lagrangian given in Eq. (3).
3.1 Pedagogical Review
We review the derivation of gravitational positivity bounds using the tree level string-state contributions to the $\phi\phi\to\phi\phi$ scattering process. Although we explicitly consider the $\phi\phi\to\phi\phi$ process as an example, the discussion is generic. We introduce two scales $\Lambda<\Lambda_{\text{QG}}$ above the electroweak scale, the new physics scale $\Lambda$ and the quantum gravity scale $\Lambda_{\text{QG}}$, and assume the following:
  • (i)
    At very high energies above the scale $\Lambda_{\text{QG}}$, gravity is UV completed by string theory. $\Lambda_{\mathrm{QG}}$ is identified with the string scale. An infinite tower of massive higher-spin states, i.e., a Regge tower, as predicted in perturbative string theory, becomes relevant for scattering amplitudes. We consider only the tree-level contributions of Regge (string) states to the scattering process.
    The scattering amplitude $\mathcal{M}(s,t)$ in the complex Mandelstam $s$-plane is assumed to satisfy
    $\lim_{|s|\to\infty}\left|\frac{\mathcal{M}(s,t<0)}{s^{2}}\right|=0,$ (6)
    where $t=0$ is excluded to avoid the $t$-channel graviton pole. This behavior in Eq. (6) is not a strong assumption, as it is satisfied by the Froissart-Martin bound [32, 43], which follows from the analyticity (i.e., causality) of the amplitude. Moreover, it also arises in tree-level string theory [56, 34].
  • (ii)
    At energies below the string scale $\Lambda_{\text{QG}}$ but above the new physics scale $\Lambda$, physics is described by a graviton coupled to a quantum field theory (QFT) describing particles and their interactions. The QFT is assumed to respect Lorentz symmetry, unitarity, and analyticity as usual. Apart from this requirement, a concrete Lagrangian is not needed. In particular, in addition to the SM and the DM particle $\phi$, various new physics contributions, i.e. particles with masses larger than $\Lambda$ and their interactions, may exist but need not be specified.
  • (iii)
    At energies below the new physics scale $\Lambda$, only the SM and the DM particle $\phi$ are present in the effective field theory (EFT). The renormalizable part of the effective Lagrangian is Eq. (3). EFT operators of dimension $n>4$, suppressed by $\Lambda^{n-4}$, arise from integrating out the heavy states (and their interactions) that are assumed to exist above $\Lambda$ (see item (ii)). Ordinary gravitational interactions, i.e., interactions with a massless graviton, are included. These are described by dimension-5, Planck-suppressed, tree-level interactions between the massless graviton and all particles via their energy-momentum tensor. Loop-level interactions of the massless graviton, which are further suppressed by the Planck scale, are neglected
We consider the 2-to-2 elastic scattering amplitude $\mathcal{M}(\phi(p_{1})\phi(p_{2})\to\phi(p_{3})\phi(p_{4}))$ and assume that the Mandelstam variable $t=(p_{3}-p_{1})^{2}$ satisfies the condition $0<-t\ll 4m_{\phi}^{2}$. The upper bound facilitates the derivation of the dispersion relations we will be using in our analysis. See Appendix A. The condition 0<t0<-t is required to avoid the $t$-channel pole arising from the massless graviton. Eventually, we take the limit $t\to-0$.
By considering the extended (almost) forward elastic scattering amplitude $\mathcal{M}(s,t)$ in the complex $s$-plane, one can relate the low-energy elastic forward amplitudes to the imaginary part of the amplitude at high energy. For details of the derivation, see Appendix A, especially the steps from Eq. (7) to Eq. (8).
The amplitude in the $s$-plane can be written as (Appendix A)
$\mathcal{M}(s,t)=\left(s-2m_{\phi}^{2}+\frac{t}{2}\right)^{2}\oint_{\mathcal{C}}\frac{ds^{\prime}}{2\pi i}\frac{\mathcal{M}(s^{\prime},t)}{(s^{\prime}-s)\left(s^{\prime}-2m_{\phi}^{2}+t/2\right)^{2}}.$ (7)
The twice-subtracted dispersion relation is (Appendix A)
$\mathcal{M}(s,t)$ $=\left[\frac{a_{-1}(t)}{s-m_{\phi}^{2}}+(s\leftrightarrow u(s,t))\right]+a_{0}+a_{1}(t)s$
$\quad+\frac{2\left(\bar{s}+\bar{t}/2\right)^{2}}{\pi}\int_{4m_{\phi}^{2}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{\left(\bar{\mu}+\bar{t}/2\right)\left[\left(\bar{\mu}+\bar{t}/2\right)^{2}-\left(\bar{s}+\bar{t}/2\right)^{2}\right]},$ (8)
where the barred variables are defined as $\bar{z}=z-4m_{\phi}^{2}/3$. The explicit forms of the functions $a_{-1}(t)$, $a_{0}$, and $a_{1}(t)$ are not relevant for our discussion here.
(a) first: $\tilde{a}_{2}(t)s^{2}$
(b) first: $\tilde{a}_{2}(t)s^{2}$
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$}hhhhGG
(c) second: $g_{2}(t)s^{2}/M^{2}_{\text{pl}}$
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$}$G$
(d) last: $Cs^{2}/(M^{2}_{\text{pl}}t)$
Figure 1: Representative contributions to the low-energy $\phi\phi\to\phi\phi$ scattering amplitude. “First,” “second,” and “last” in the subcaptions refer to the corresponding terms in Eq. (9).
From assumption (iii), the terms proportional to $s^{2}$ in the low-energy amplitude ($E<\Lambda$) are
$\mathcal{M}(s,t)\supset a_{2}(t)s^{2}=\tilde{a}_{2}(t)s^{2}-g_{2}(t)s^{2}/M^{2}_{\text{pl}}-C\frac{s^{2}}{M^{2}_{\text{pl}}t},$ (9)
where we neglect terms not proportional to $s^{2}$. (Including them will restore crossing symmetry.) As shown in Figure 1, the contributions on the right-hand side of Eq. (9) originate from four types of diagrams (a)–(d):
  • (a)
    The first term receives loop-level contributions from the Higgs-portal DM.
  • (b)
    The first term also receives tree-level contributions from dimension-8 derivative operators, such as $(\partial\phi)^{4}/\Lambda^{4}$.
  • (c)
    The second term corresponds to a $t$-channel graviton contribution with a loop of SM and $\phi$ particles.
  • (d)
    The last term arises from a tree-level $t$-channel graviton exchange, with a positive constant $C$ and the Planck mass $M_{\text{pl}}$.
For cases (a) and (c), the full contributions of diagrams involving the Higgs-portal DM are discussed in Sec. 3.2. The last term, i.e., the $t$-channel contribution from a massless graviton, $\mathcal{M}(s,t)\sim-Cs^{2}/(M^{2}_{\text{pl}}t)$, diverges in the forward limit ($t\to-0$). Using $u=4m_{\phi}^{2}-s-t$ and taking of the limit $t\to-0$ at the end of the derivation, only the Mandelstam $s$ contribution remains.
A term proportional to $s^{2}$ arises only from the final term of the dispersion relation in Eq. (8):
$\mathcal{M}(s,t)\supset 2\frac{\left(\bar{s}+\bar{t}/2\right)^{2}}{\pi}\int_{4m_{\phi}^{2}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\bar{\mu}+\bar{t}/2)\left[(\bar{\mu}+\bar{t}/2)^{2}-(\bar{s}+\bar{t}/2)^{2}\right]}.$ (10)
From Eqs. (9) and (10), by taking the second derivative of the amplitude with respect to $s$ at $s=2m_{\phi}^{2}-t/2$, we obtain
$2\tilde{a}_{2}(t)-2g_{2}(t)/M^{2}_{\text{pl}}-2\frac{C}{M^{2}_{\text{pl}}t}$ $=\frac{4}{\pi}\int_{4m_{\phi}^{2}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu-2m_{\phi}^{2}+t/2)^{3}}.$ (11)
Here, we include the $u$-channel contributions in the LHS of Eq. (11) via $u=4m_{\phi}^{2}-s-t\sim 4m_{\phi}^{2}-s$, which are absorbed into $\tilde{a}_{2}(t)$ and $g_{2}(t)$, respectively. The right-hand side of Eq. (11) can be divided into three energy regions:
[…]
Following the analysis of Ref. [17, 28, 29] we rearrange terms of Eq. (11) into the form
$2\tilde{a}_{2}(t)-\frac{4}{\pi}\int_{4m_{\phi}^{2}}^{\Lambda^{2}}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu-2m_{\phi}^{2}+t/2)^{3}}-2g_{2}(t)/M^{2}_{\text{pl}}$ (13)
$=$ $2\frac{C}{M^{2}_{\text{pl}}t}+\frac{4}{\pi}\int_{\Lambda_{\mathrm{QG}}^{2}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu-2m_{\phi}^{2}+t/2)^{3}}+\frac{4}{\pi}\int_{\Lambda^{2}}^{\Lambda_{\mathrm{QG}}^{2}}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu-2m_{\phi}^{2}+t/2)^{3}}.\vphantom{\Bigg|}$ (14)
In the forward limit $t\to-0$, the first term on the right-hand side of Eq. (14) diverges to -$\infty$ since $C>0$. However, this divergence is canceled by the contributions of tree-level Regge states in the first integral, corresponding to energies $E>\Lambda_{\mathrm{QG}}$. Above the string or quantum gravity scale, $\Lambda_{\mathrm{QG}}=\alpha^{\prime-1/2}$, a tower of massive higher spin states (a Regge tower) appear, as assumed in assumption (i). The scattering amplitude receives contributions from these Regge states [56]. Here, the scale $\alpha^{\prime}$ corresponds to the usual $\alpha^{\prime}$ parameter in string theory. The associated divergence is almost canceled by summing the contributions of the Regge states, leaving at most a possible remnant of order $\mathcal{O}(\alpha^{\prime}/M^{2}_{\mathrm{pl}})=\mathcal{O}(1/(\Lambda_{\mathrm{QG}}^{2}M^{2}_{\mathrm{pl}}))$, as shown by Tokuda et al. [56] using string theory. (See also Refs. [13, 34].)
Now, Eq. (14) can be rewritten by neglecting all string states except for the Regge states and by omitting loop-level contributions in the UV region, $E>\Lambda_{\mathrm{QG}}$:
$2\tilde{a}_{2}(t\to-0)-\frac{4}{\pi}\int_{4m_{\phi}^{2}}^{\Lambda^{2}}d\mu\frac{\mathrm{Im}\,\mathcal{M}(\mu+i\epsilon,t\to-0)}{(\mu-2m_{\phi}^{2})^{3}}-\frac{2g_{2}(t\to-0)}{M_{\mathrm{pl}}^{2}}$
$\qquad=\frac{4}{\pi}\int_{\Lambda^{2}}^{\Lambda_{\mathrm{QG}}^{2}}d\mu\frac{\mathrm{Im}\,\mathcal{M}(\mu+i\epsilon,t\to-0)}{(\mu-2m_{\phi}^{2})^{3}}>0.$ (15)
Here, we take the forward limit $t\to-0$. Note that the right-hand side of Eq. (15) is manifestly positive due to unitarity, assumption (ii), since unitarity leads to the optical theorem for forward 2-to-2 elastic scattering:
$\mathrm{Im}\,\mathcal{M}(p_{1},p_{2}\to p_{1},p_{2})=s\,\sigma_{\mathrm{tot}}(p_{1},p_{2}\to\text{anything})>0\quad(\text{in the massless limit}).$ (16)
Note that the denominator of the integrand on the right-hand side of Eq. (15) is also positive as $\Lambda\gg m_{\phi}$ at the lower limit of the integral. This implies that the left-hand side of Eq. (15) is positive, despite the second and third terms contributing with minus signs.
Therefore, Eq. (15) can be written as
$B(\Lambda)$ $=B_{\text{non-grav}}+B_{\text{grav}}>0,$ (17)
$B_{\text{non-grav}}$ $=2\tilde{a}_{2}(t\to-0)-\frac{4}{\pi}\int_{4m^{2}_{\phi}}^{\Lambda^{2}}d\mu\frac{\mathrm{Im}\,\mathcal{M}(\mu+i\epsilon,t\to-0)}{(\mu-2m^{2}_{\phi})^{3}},$ (18)
$B_{\text{grav}}$ $=-\frac{2g_{2}(t\to-0)}{M^{2}_{\text{pl}}}=\lim_{t\to-0}\left(\frac{\partial^{2}\mathcal{M}_{\text{grav}}(s,t)}{\partial s^{2}}\right)_{s=2m^{2}_{\phi}}.$ (19)
Here, $B_{\text{grav}}$ represents the subtraction of the $s^{2}$ dependence of the amplitude, so its mass dimension is 4-4. Non-gravitational part, $B_{\text{non-grav}}$, likewise has mass dimension 4-4, and it involves the subtraction of a positive contribution from the integrand.
Neglecting the graviton contributions (i.e., terms suppressed by $1/M^{2}_{\text{pl}}$, corresponding to $B_{\text{grav}}$) and using Eq. (14), $B_{\text{non-grav}}$ can be expressed by ignoring the UV details of the string states, as
$B_{\text{non-grav}}=\frac{4}{\pi}\int_{\Lambda^{2}}^{\infty}d\mu\frac{\mathrm{Im}\,\mathcal{M}_{\mathrm{w/o~Regge}}(\mu+i\epsilon,t\to-0)}{(\mu-2m^{2}_{\phi})^{3}}>0.$ (20)
We will use Eq. (20) to estimate $B_{\text{non-grav}}$ in Sec. 3.2.
In Eq. (17), $B_{\text{grav}}$ arises from the contribution involving a $t$-channel massless graviton (see panel (c) in Figure 1). In many cases, $B_{\text{grav}}$ is negative [24, 23, 10, 4, 12, 48]. Consequently, the gravitational positivity bound in Eq. (17) reads
$B_{\text{non-grav}}>|B_{\text{grav}}|.$ (21)
Further discussions on the origin and validity of these bounds through string-loop effects in string-inspired models can be found in Refs. [5, 6, 36, 27, 34, 21]. Related insights from the Weak Gravity Conjecture appear in Refs. [35, 10, 22, 16, 46, 1].
Thus, Eqs. (20) and (21) show that, if we assume only renormalizable interactions and no additional new physics above the scale $E>\Lambda$, then a violation of the positivity condition in Eq. (21) at $\Lambda=E$ would imply that new physics at or above $E$ is required to raise the cutoff energy scale $\Lambda$ above $E$, since such new states must provide additional positive contributions to the left-hand side, $B_{\text{non-grav}}$. We adopt this assumption in Sec. 3.2 to determine the energy scale up to which the renormalizable Higgs-portal DM remains valid.
Finally, let us comment on the conventional argument for positivity bounds. Dimension-8 EFT operators are included in $\tilde{a}_{2}(t\to-0)$ in Eq. (18), in addition to the Higgs-portal DM contributions. If new physics exists only at energy scales above $\Lambda$, then Eqs. (18) and (20) imply,
$\tilde{a}_{2}(t\to-0)=\frac{4}{\pi}\int_{\Lambda^{2}}^{\infty}d\mu\frac{\mathrm{Im}\,\mathcal{M}_{\mathrm{w/o~Regge}}(\mu+i\epsilon,t\to-0)}{(\mu-2m^{2}_{\phi})^{3}}>0.$ (22)
Here, $\tilde{a}_{2}(t\to-0)$ corresponds to a combination of the Wilson coefficients of dimension-8 operators (for example, see [57]). Eq. (22) shows that these dimension-8 contributions on the LHS arise from heavy states appearing at or above the energy scale $\Lambda$. By unitarity of the new physics sector, these states provide positive contributions to the RHS via $\mathrm{Im}\,\mathcal{M}$. This is the standard argument for the original positivity bounds [3].
3.2 Gravitational Positivity Bounds with $\phi\phi\to\phi\phi$
Figure 2: Non-gravitational Feynman diagrams for $\phi\phi\to\phi\phi$
We now compute $B(\Lambda)$ for the $\phi\phi\to\phi\phi$ scattering process at the one-loop level, following Sec. 3.1. The one-loop amplitudes are calculated using FeynRules [25, 7] and the Mathematica packages FeynArts [33], FeynCalc [45, 52, 53, 54], and Package-X [50].
We first present the results for the non-gravitational contributions to the $\phi\phi\to\phi\phi$ scattering process. Figure 2 shows the leading non-gravitational one-loop Feynman diagrams for this process.
As illustrated there, only the Higgs-portal contributions are included, whereas possible contributions from new physics, i.e., heavy states, are not considered. (We do not include EFT operators in our calculation.) Consequently, if our positivity bounds are violated at a certain energy scale $\Lambda=E$, additional new physics beyond the Higgs-portal interactions would be required to raise the cutoff energy scale $\Lambda$ above $E$, as discussed in Sec. 3.1. If $(\Lambda=)E$ is sufficiently high, e.g., the grand unified theory scale or a typical string scale, this implies that the renormalizable Higgs-portal DM framework in Eq. (1) (without any additional new physics) remains valid up to that high energy scale from the perspective of positivity.
The candy-like diagrams in Figure 2 contribute to $B_{\text{non-grav}}(\Lambda)$ with a suppression of order $1/\Lambda^{4}$, whereas other diagrams are further suppressed, e.g., by factors of $v^{2}/\Lambda^{6}$ or $v^{4}/\Lambda^{8}$. The dominant non-gravitational contribution, $B_{\text{non-grav}}(\Lambda)$, is therefore given by
$B^{\phi\phi\to\phi\phi}_{\text{non-grav}}$ $=\frac{\lambda_{h\phi}^{2}+\lambda_{\phi}^{2}}{16\pi^{2}\Lambda^{4}}.$ (23)
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$}hhhhGG
[…]
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi$}hhhhGG
[…]
Figure 3: $t$-channel graviton exchange Feynman diagrams for $\phi\phi\to\phi\phi$
Next, we present the gravitational contribution, $B_{\text{grav}}(\Lambda)$, for the $\phi\phi\to\phi\phi$ process. Figure 3 shows the leading one-loop gravitational Feynman diagrams in the $t$-channel. The dominant gravitational contribution is given by
$B^{\phi\phi\to\phi\phi}_{\text{grav}}$ $=B^{\phi\phi\to\phi\phi}_{\text{grav(a)}}+B^{\phi\phi\to\phi\phi}_{\text{grav(b)}}+B^{\phi\phi\to\phi\phi}_{\text{grav(c)}}+B^{\phi\phi\to\phi\phi}_{\text{grav(d)}}=-\frac{\lambda^{2}_{h\phi}v^{2}}{24\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h}),$ (24)
$f(x)$ $=\frac{1}{x^{6}(1-4x^{2})^{2}}\biggl\{28x^{6}-15x^{4}+2x^{2}-(1-4x^{2})^{2}(x^{2}-1)\ln(x^{2})$
$\,\,\,\,\,\,-2\sqrt{1-4x^{2}}(2x^{6}-12x^{4}+7x^{2}-1)\ln\left[\frac{1}{2x}\left(\sqrt{1-4x^{2}}+1\right)\right]\biggr\},$ (25)
where $\overline{M}_{\text{pl}}=M_{\text{pl}}/\sqrt{8\pi}\sim 2.4\times 10^{18}~\mathrm{GeV}$ is the reduced Planck mass, and $m_{h}$ is the Higgs boson mass. Explicit expressions for each contribution, i.e., $B^{\phi\phi\to\phi\phi}_{\text{grav(a/b/c/d)}}$, are provided in Appendix B.
Figure 4: Plots of $f(x)$ (left) and $f^{-1/4}(x)$ (right). The function $f(x)$ is real and positive in the plotted range.
The function $f(x)$ is displayed in the left panel of Figure 4. As can be seen from the plot, this function is real and positive. This implies that the gravitational contribution in Eq. (24) is negative.
Combining Eqs. (23) and (24), the gravitational positivity bound for the $\phi\phi\to\phi\phi$ process is
$B^{\phi\phi\to\phi\phi}(\Lambda)$ $=B^{\phi\phi\to\phi\phi}_{\text{non-grav}}+B^{\phi\phi\to\phi\phi}_{\text{grav}}\geq 0,$ (26)
$\lambda_{h\phi}^{2}+\lambda_{\phi}^{2}$ $\geq\frac{2\lambda^{2}_{h\phi}\Lambda^{4}v^{2}}{3\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f(m_{\phi}/m_{h}),$ (27)
which explicitly gives
$\Lambda$ $\leq\left(\frac{3}{2}\right)^{1/4}\left(\frac{\overline{M}_{\text{pl}}}{v}\right)^{1/2}\left(1+\frac{\lambda^{2}_{\phi}}{\lambda^{2}_{h\phi}}\right)^{1/4}f^{-1/4}(m_{\phi}/m_{h})m_{h}\,,\quad\text{or}\quad\lambda_{h\phi}=0.$ (28)
From Eq. (28), when $\lambda_{h\phi}\neq 0$ and $\lambda_{\phi}=0$, i.e., in the presence of the Higgs-portal interaction without scalar DM self-interactions, we obtain
$\Lambda$ $\leq\left(\frac{3}{2}\right)^{1/4}\left(\frac{\overline{M}_{\text{pl}}}{v}\right)^{1/2}f^{-1/4}(m_{\phi}/m_{h})m_{h},$ (29)
$\Lambda$ $\lesssim 10^{10}~\mathrm{GeV}.$ (30)
From Eq. (30), new physics is expected to appear below approximately $10^{10}$ GeV when the Higgs-portal scalar DM has no self-interaction. Here, we assume that $f^{-1/4}(m_{\phi}/m_{h})$ is of order $\mathcal{O}(1)$. Note that $f(x)\to 1/3$ as $x\to 0$, corresponding to $f^{-1/4}(x)\sim 1.3$. Therefore, we focus on the light DM region. The function $f^{-1/4}(x)$, shown in the right panel of Figure 4, takes values of order $\mathcal{O}(1)$ when $m_{\phi}<m_{h}$, supporting our focus on this DM mass region. As the figure suggests, heavier DM masses may allow $\Lambda$ to reach the GUT scale. Indeed, this occurs if the DM mass exceeds approximately $10^{10}$ GeV. However, the light DM region with $m_{\phi}<m_{h}$ is more interesting to investigate, as it is more severely constrained by the gravitational positivity bounds.
Figure 5: Parameter space for $\lambda_{\phi}/\lambda_{h\phi}$ versus $\Lambda$ for light DM with $m_{\phi}<m_{h}$, satisfying the gravitational positivity bound from the $\phi\phi\to\phi\phi$ process. The green region indicates where the bound in Eq. (28) is satisfied.
We illustrated the parameter space of $\lambda_{\phi}/\lambda_{h\phi}$ versus $\Lambda$ in Figure 5 for non-zero $\lambda_{\phi}$. The gravitational positivity bounds in Eq. (28) are satisfied in the green region. As expected from Eq. (28), comparing the cases $\lambda_{\phi}=0$ and $\lambda_{\phi}/\lambda_{h\phi}=100$ in Figure 5, the upper bound on the new physics scale $\Lambda$ increases by approximately one order of magnitude. If there is a large hierarchy between the two couplings, such that the Higgs-portal coupling is small (i.e., $\lambda_{\phi}/\lambda_{h\phi}=10^{2n}$ with $n=1,2,\dots$), the upper bound on the new physics scale is approximately given by
$\Lambda$ $\lesssim 10^{10+n}~\mathrm{GeV},$
$\lambda_{\phi}/\lambda_{h\phi}$ $=10^{2n}\quad(n=1,2,\cdots),$
$\text{for }\,m_{\phi}$ $<m_{h}.$ (31)
4 Phenomenological Implications for Light Dark Matter
Based on the results in Sec. 3, we discuss the implications of the gravitational positivity bounds for the Higgs-portal light real scalar DM model. From Eq. (31), if the new physics scale $\Lambda$ corresponds to the GUT scale or a typical string scale, i.e. $\Lambda\sim 10^{16}$ GeV, the bound implies a large hierarchy between the Higgs-portal coupling and the DM self-coupling:
$\lambda_{\phi}/\lambda_{h\phi}$ $=10^{12}\quad(\text{for $\Lambda\sim$ GUT/typical string scale}),$
$\text{for }m_{\phi}$ $<m_{h}.$ (32)
From Eq. (32) and requiring perturbativity, i.e., $\lambda_{\phi}\lesssim 4\pi$, it follows that $\lambda_{h\phi}\ll 1$. This implies that the Higgs-portal coupling must be extremely small, if it exists at all. When $\lambda_{h\phi}\ll 1$ but non-zero, the Higgs-portal DM freeze-in scenario becomes relevant [19, 41].
We consider two cases for light DM ($m_{\phi}<m_{h}$): $\lambda_{\phi}\lesssim 10^{-2}$ and $\lambda_{\phi}\gtrsim 10^{-2}$. The latter case is necessary to realize the GUT-scale or typical string-scale $\Lambda$.
  • Case 1: ($\lambda_{\phi}\lesssim 10^{-2}$)
    For light DM, i.e., $m_{\phi}<m_{h}$, the process $h\to\phi\phi$ dominates over $hh\to\phi\phi$ [41] in determining the DM relic abundance. For 100 MeV DM [19],
    $\lambda_{\phi}/\lambda_{h\phi}\lesssim 10^{10}.$ (33)
    This corresponds to $\Lambda\lesssim 10^{15}$ GeV according to Eq. (31). Since a smaller DM mass lowers $\Lambda$, this scenario does not lead to a GUT-scale $\Lambda$.
  • Case 2: ($\lambda_{\phi}\gtrsim 10^{-2}$)
    Here, the DM self-interaction plays a significant role in DM production via dark freeze-out [19, 31, 2]. With a non-zero Higgs-portal coupling, the following parameter set can simultaneously realize both the GUT-scale $\Lambda$ (or typical string scale $\Lambda$) and the correct DM relic abundance (see Fig. 5 in [19]):
    $\lambda_{h\phi}$ $\sim 2\times 10^{-12},$
    $\lambda_{\phi}$ $\sim 2,$
    $m_{\phi}$ $\sim 300~\mathrm{MeV}.$ (34)
    Additionally, this parameter set (34) satisfies stringent astrophysical constraints from Milky Way satellite dwarf galaxies, i.e., $\sigma/m<0.2~\mathrm{cm}^{2}/\mathrm{g}$ [9], where $\sigma/m$ is the DM self-interaction cross section divided by its mass.
    The dark freeze-out scenario proceeds as follows [19]. At temperatures below the Higgs mass (but still relatively high), Higgs decays ($h\to\phi\phi$) produce an initial population of DM particles. Once a sufficient number of DM particles is produced, $2\to 4$ DM self-interactions rapidly increase the DM number density. After chemical equilibrium is reached, the DM number density tracks the equilibrium value until the particles become non-relativistic. Eventually, as the temperature $T$ (or dark temperature $T_{D}$) decreases, freeze-out occurs when $m_{\phi}/T\sim 0.3$ (or $m_{\phi}/T_{D}\sim 7$), fixing the DM yield.
In conclusion, the gravitational positivity bound for light DM, i.e., $m_{\phi}<m_{h}$, indicates that realizing the GUT-scale or typical string-scale $\Lambda$ in the Higgs-portal model requires a large hierarchy between an $\mathcal{O}(1)$ DM self-coupling and an extremely small Higgs-portal coupling of $\mathcal{O}(10^{-12})$, together with a sub-GeV DM mass.
For the parameter set given in Eq. (34), and with $\Lambda\lesssim 10^{16}$ GeV, the possible additional contribution on the RHS of Eq. (17), which could potentially lead to a positivity violation, is of order $\mathcal{O}\bigl(1/(\Lambda^{2}_{\mathrm{QG}}M^{2}_{\text{pl}})\bigr)=\mathcal{O}(\alpha^{\prime}/M^{2}_{\text{pl}})$ on the RHS of Eq. (17) can be safely neglected. Here, we assume a typical value of $\Lambda_{\mathrm{QG}}=\alpha^{\prime-1/2}\sim 10^{16}$ GeV, representing the scale of Reggeization [35, 56, 12, 48].
Finally, we comment on the gravitational contribution $B_{\mathrm{grav}}$ associated with the four-point self-interaction of DM. In this case, the one-loop contribution is absent, and the leading contribution arises at the two-loop level [48]. As a result, $B_{\mathrm{grav}}$ itself, which provides a lower bound, receives an additional contribution scaling as $\mathcal{O}\!\left(\lambda_{\phi}^{2}/(m^{2}_{\phi}\overline{M}^{2}_{\mathrm{pl}})\right)$ [48]. For sub-GeV DM with a GUT-scale (or lower) cutoff, however, this contribution can be safely neglected in comparison with the non-gravitational contribution from the self-coupling, which scales as $\sim\lambda_{\phi}^{2}/\Lambda^{4}$ in $B_{\mathrm{non\text{-}grav}}$. For DM masses lighter than the sub-GeV scale, the positivity bounds become more stringent, but this does not alter our conclusion, as our bound is conservative. For heavier DM masses, the two-loop gravitational contribution is further suppressed and can therefore be neglected.
5 Summary and Discussion
We have considered a Higgs-portal real scalar DM model with a $Z_{2}$-odd parity and renormalizable interactions. In this framework, we applied gravitational positivity bounds to the forward DM scattering process $\phi\phi\to\phi\phi$. For light DM with $m_{\phi}<m_{h}$ and in the absence of a DM self-coupling, new physics beyond the Higgs-portal scenario is required to appear below an energy scale of $10^{10}$ GeV.
When a DM self-coupling is present, a large hierarchy between the self-coupling $\lambda_{\phi}$ and the Higgs-portal coupling $\lambda_{h\phi}$ raises the upper bound on the new physics scale $\Lambda$. For example, with $\lambda_{\phi}/\lambda_{h\phi}=10^{12}$, the gravitational positivity bound allows a $\Lambda$ at the GUT scale or at a typical string scale. In such a scenario, the dark freeze-out mechanism can still account for the observed relic abundance of DM in the Universe. We find that parameter values of $\lambda_{\phi}\sim\mathcal{O}(1)$, $\lambda_{h\phi}\sim 10^{-12}$, and sub-GeV DM are consistent with achieving a GUT-scale or string-scale $\Lambda$ within the Higgs-portal light DM framework.
We comment on the possibility of obtaining additional constraints from gravitational positivity. Processes suitable for positivity bounds must involve stable particles [14, 11]. For the process $\phi e\to\phi e$, infrared (IR) divergences in $B_{\text{grav}}$ pose a significant challenge. While redefining asymptotic electron states using the Faddeev-Kulish formalism can resolve IR divergences in QED [49], introducing a $t$-channel graviton exchange generates new contributions from diagrams with two photon propagators (and one electron propagator) in the triangle loop, leading to even more severe IR divergences from photons.
In the case of the process $\phi\gamma\to\phi\gamma$, the $B_{\text{non-grav}}$ contribution appears first at the two-loop level. In principle, this contribution can be computed via the optical theorem, providing a potential pathway for further refinement of the bounds.
Acknowledgements
We have used the package TikZ-Feynman [30] to draw the Feynman diagrams. We would like to express our sincere gratitude to Katsuki Aoki for useful discussions and comments, Kaoru Hagiwara for his encouragement, Akio Sugamoto for his valuable input on infrared divergences, Tatsu Takeuchi for carefully reading the manuscript and providing valuable feedback, and Kazuya Yonekura for his helpful comments on string theory. We also thank the Astrophysics & Cosmology Group at the Yukawa Institute for Theoretical Physics, Kyoto University, for their warm hospitality, where a part of this work was carried out. This work is supported by JSPS KAKENHI Grant Number JP24K17040.
Appendix A Derivation from Eq. (7) to Eq. (8)
Figure 6: Scattering amplitude $\mathcal{M}(s,t)$ in the complex $s$-plane with poles and branch cuts. The integration contour $\mathcal{C}$ in the upper-left panel can be deformed into the contours $\mathcal{C}^{\prime}$ and $\mathcal{C}_{\infty}$ in the upper-right panel, following the analytic structure of the amplitude aside from poles and branch cuts. The bottom-left panel shows how this deformation is achieved by introducing two paths, $A$ and $A^{\prime}$, which can be brought arbitrarily close to each other and ultimately removed.
From assumption (ii), valid above the energy scale $E>\Lambda$, the scattering amplitude obeys Lorentz invariance, analyticity (except for poles and branch cuts), crossing symmetry (from the LSZ reduction formula), and unitarity. Using this, we employ a twice-subtracted dispersion relation, Eq. (8), proportional to $\mathrm{Im}\,\mathcal{M}(s,t)/s^{3}$, to relate the real part of the amplitude, Eq. (9). This is justified by the high-energy behavior in Eq. (6) (assumption (i)) and is used to extract the leading contribution to the amplitude relevant for the gravitational positivity bounds.
Let us consider the (nearly) forward elastic amplitude $\mathcal{M}(s,t)$ in the complex $s$-plane, as shown in Figure 6. The contour $\mathcal{C}$ in Eq. (7) can be deformed into $\mathcal{C}^{\prime}$ and $\mathcal{C}_{\infty}$.
With the contour $\mathcal{C}$ shown in the upper-left panel of Figure 6, the amplitude can be written as
$\mathcal{M}(s,t)=\left(s-2m_{\phi}^{2}+\frac{t}{2}\right)^{2}\oint_{\mathcal{C}}\frac{ds^{\prime}}{2\pi i}\frac{\mathcal{M}(s^{\prime},t)}{(s^{\prime}-s)\left(s^{\prime}-2m_{\phi}^{2}+t/2\right)^{2}},$ (7)
where the factor $\left(s-2m_{\phi}^{2}+t/2\right)^{2}$ [13] ensures that the $u$-channel amplitude, arising from crossing symmetry, takes the same functional form.
Deforming $\mathcal{C}$ into $\mathcal{C}^{\prime}$ and $\mathcal{C}_{\infty}$ as in Figure 6 leads to the twice-subtracted dispersion relation,
$\mathcal{M}(s,t)$ $=\left[\frac{a_{-1}(t)}{s-m_{\phi}^{2}}+(s\leftrightarrow u(s,t))\right]+a_{0}+a_{1}(t)s$
$\quad+2\frac{\bigl(\bar{s}+\bar{t}/2\bigr)^{2}}{\pi}\int_{4m_{\phi}^{2}}^{\infty}d\mu\,\frac{\mathrm{Im}\,\mathcal{M}(\mu+i\epsilon,t)}{\bigl(\bar{\mu}+\bar{t}/2\bigr)\bigl[(\bar{\mu}+\bar{t}/2)^{2}-(\bar{s}+\bar{t}/2)^{2}\bigr]},$ (8)
where barred variables are defined by $\bar{z}=z-4m_{\phi}^{2}/3$.
The explicit forms of $a_{-1}(t)$, $a_{0}$, and $a_{1}(t)$ are not required for our analysis. The bottom-left panel of Figure 6 illustrates how the upper-right panel is obtained by introducing two nearly coincident paths, $A$ and $A^{\prime}$, which can eventually be shrunk to zero. The contribution from $\mathcal{C}_{\infty}$ vanishes because of assumption (i) in the limit $|s|\to\infty$, as shown later in this appendix.
To we outline the derivation of Eq. (8) from Eq. (7), let us first consider a part of the contour $\mathcal{C}^{\prime}$ contribution. The sus-u crossing symmetry for the scalar field is used as
$\mathcal{M}(s,t)=\mathcal{M}(u,t).$ (35)
With the on-shell condition $u=4m^{2}_{\phi}-s-t$,
$\mathcal{M}(s,t)=\mathcal{M}(4m^{2}_{\phi}-s-t,t).$ (36)
This implies that for $(0<)-t\ll 4m^{2}_{\phi}$, when the amplitude has a pole at $s=m^{2}_{\phi}$, it also has a pole at $s\sim 3m^{2}_{\phi}$. Both originate from the tree-level forward elastic scattering $\phi\phi\to\phi\phi$ mediated by $\phi$ in the $s$- and $u$-channels. By the residue theorem,
$\oint_{\mathcal{C}^{\prime}}\frac{ds^{\prime}}{2\pi i}\frac{\mathcal{M}(s^{\prime},t)}{(s^{\prime}-s)\left(s^{\prime}-2m^{2}_{\phi}+t/2\right)^{2}}$ $=-\frac{\mathrm{Res}_{s^{\prime}=m^{2}_{\phi}}\mathcal{M}(s^{\prime},t)}{(m^{2}_{\phi}-s)\left(-m^{2}_{\phi}+t/2\right)^{2}}+(\mathrm{crossing}).$ (37)
To obtain the relevant part of Eq. (7), we multiply $(s-2m^{2}_{\phi}+t/2)^{2}$ to the above Eq. (37). Here,
$(s-2m^{2}_{\phi}+t/2)^{2}=(-m^{2}_{\phi}+t/2-m^{2}_{\phi}+s)^{2}=(m^{2}_{\phi}-t/2)^{2}(1-(s-m^{2}_{\phi})/(m^{2}_{\phi}-t/2))^{2}.$ (38)
Thus, multiplying Eq. (37) by the factor $(s-2m^{2}_{\phi}+t/2)^{2}$ gives
$\left(s-2m^{2}_{\phi}+\frac{t}{2}\right)^{2}\oint_{\mathcal{C}^{\prime}}\frac{ds^{\prime}}{2\pi i}\frac{\mathcal{M}(s^{\prime},t)}{(s^{\prime}-s)\left(s^{\prime}-2m^{2}_{\phi}+t/2\right)^{2}}=\left[\frac{a_{-1}(t)}{s-m^{2}_{\phi}}+(s\leftrightarrow u(s,t))\right]+a_{0}+a_{1}(t)s,$ (39)
with functions $a_{-1}(t)$, $a_{0}$, and $a_{1}(t)$. Eq. (39) reproduces the first four terms in Eq. (8), corresponding to pole contributions.
Next, we see the contribution from $\mathcal{C}_{\infty}$ vanishes.
$\left|\oint_{\mathcal{C}_{\infty}}ds^{\prime}\frac{\mathcal{M}(s^{\prime},t)}{(s^{\prime}-s)(s^{\prime}-2m^{2}_{\phi}+t/2)^{2}}\right|$ $=\left|\oint_{\mathcal{C}_{\infty}}ds^{\prime}\frac{\mathcal{M}(s^{\prime},t)}{s^{\prime 3}}\right|$
$\sim\lim_{|s^{\prime}|\to\infty}\left|\frac{\mathcal{M}(s^{\prime},t)}{s^{\prime 2}}\right|=0,$ (40)
where the last step uses Eq. (6) in (i) with $t<0$. The reason for dividing $\mathcal{M}(s,t)$ by $s^{3}$ is precisely to suppress this contribution and make the dispersion relation well-defined. By doing so, we can ignore the $\mathcal{C}_{\infty}$ contribution.
The remaining contribution arises from the branch cuts in the complex s-plane. These appear when s reaches the on-shell energy of two $\phi$ particles, i.e. from one-loop intermediate states that generate an imaginary part of the amplitude. Explicitly, the relevant pieces are
$\left(\int_{4m^{2}_{\phi}}^{\infty}+\int_{-\infty}^{0}\right)d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$+\left(\int_{\infty}^{4m^{2}_{\phi}}+\int_{0}^{-\infty}\right)d\mu\frac{\mathcal{M}(\mu-i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}.$ (41)
Because of the on-shell condition $s+t+u=4m^{2}_{\phi}$ with $t\ll 4m^{2}_{\phi}$, the branch cut for $s\geq 4m^{2}_{\phi}$ corresponds to the cut for $u\lesssim 0$. By crossing symmetry, $\mathcal{M}(s,t)=\mathcal{M}(u,t)$, another branch cut starts from $s=0$ to the negative $s$. We introduce the shifted variables $\bar{z}=z-(4m^{2}_{\phi}/3)$, so that $\bar{s}+\bar{t}+\bar{u}=0$ from the on-shell condition.
The positive real $s$ parts of Eq. (41) become
$\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}+\int_{\infty}^{4m^{2}_{\phi}}d\mu\frac{\mathcal{M}(\mu-i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$\,\,\,\,=\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}-\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathcal{M}(\mu-i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$\,\,\,\,=\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}-\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathcal{M}^{*}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$\,\,\,\,=2i\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}.$ (42)
We use the Schwarz reflection principle $\mathcal{M}(s^{*},t)=\mathcal{M}^{*}(s,t)$ from the Analyticity of the amplitude from the second to the third lines. Similarly, the negative-real-$s$ part in Eq. (41) is
$\int_{-\infty}^{0}d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}+\int_{0}^{-\infty}d\mu\frac{\mathcal{M}(\mu-i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$\,\,\,\,=-\int_{0}^{-\infty}d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu-s)(\mu-2m^{2}_{\phi}+t/2)^{2}}+\int_{0}^{\infty}d\mu\frac{\mathcal{M}(-\mu-i\epsilon,t)}{(\mu+s)(\mu+2m^{2}_{\phi}-t/2)^{2}}$
$\,\,\,\,=-\int_{0}^{\infty}d\mu\frac{\mathcal{M}(-\mu+i\epsilon,t)}{(\mu+s)(\mu+2m^{2}_{\phi}-t/2)^{2}}+\int_{0}^{\infty}d\mu\frac{\mathcal{M}(-\mu-i\epsilon,t)}{(\mu+s)(\mu+2m^{2}_{\phi}-t/2)^{2}}$
$\,\,\,\,=-\int_{0}^{\infty}d\mu\frac{\mathcal{M}(4m^{2}_{\phi}+\mu-i\epsilon-t,t)}{(\mu+s)(\mu+2m^{2}_{\phi}-t/2)^{2}}+\int_{0}^{\infty}d\mu\frac{\mathcal{M}(4m^{2}_{\phi}+\mu+i\epsilon-t,t)}{(\mu+s)(\mu+2m^{2}_{\phi}-t/2)^{2}}$
$\,\,\,\,=\int_{4m^{2}_{\phi}-t}^{\infty}d\mu\frac{\mathcal{M}(\mu+i\epsilon,t)}{(\mu+t-4m^{2}_{\phi}+s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$\,\,\,\,\,\,\,\,-\int_{4m^{2}_{\phi}-t}^{\infty}d\mu\frac{\mathcal{M}(\mu-i\epsilon,t)}{(\mu+t-4m^{2}_{\phi}+s)(\mu-2m^{2}_{\phi}+t/2)^{2}}$
$\,\,\,\,=2i\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu+t-4m^{2}_{\phi}+s)(\mu-2m^{2}_{\phi}+t/2)^{2}}.$ (43)
Here we employed crossing symmetry and the on-shell condition, e.g., $\mathcal{M}(-s+i\epsilon)=\mathcal{M}(4m^{2}_{\phi}-(-s+i\epsilon)-t)=\mathcal{M}(4m^{2}_{\phi}+s-i\epsilon-t)$ and $\mathcal{M}(-s-i\epsilon)=\mathcal{M}(4m^{2}_{\phi}-(-s-i\epsilon)-t)=\mathcal{M}(4m^{2}_{\phi}+s+i\epsilon-t)$. Additionally, we changed the integration variable from $\mu$ to $\mu-4m^{2}_{\phi}+t$, used the condition of $t\ll 4m^{2}_{\phi}$, and applied the Schwarz reflection principle.
Adding Eqs. (42) and (43) gives
$2i\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{(\mu-2m^{2}_{\phi}+t/2)^{2}}\left(\frac{1}{\mu-s}+\frac{1}{(\mu+t-4m^{2}_{\phi}+s)}\right)$
$\,\,\,\,=4i\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{\left(\bar{\mu}+\bar{t}/2\right)\left[\left(\bar{\mu}+\bar{t}/2\right)^{2}-\left(\bar{s}+\bar{t}/2\right)^{2}\right]}.$ (44)
By multiplying Eq. (44) by $(s-2m_{\phi}^{2}+t/2)^{2}=(\bar{s}+\bar{t}/2)^{2}$ and dividing it by $2\pi i$, we obtain
$2\frac{\left(\bar{s}+\bar{t}/2\right)^{2}}{\pi}\int_{4m^{2}_{\phi}}^{\infty}d\mu\frac{\mathrm{Im}\mathcal{M}(\mu+i\epsilon,t)}{\left(\bar{\mu}+\bar{t}/2\right)\left[\left(\bar{\mu}+\bar{t}/2\right)^{2}-\left(\bar{s}+\bar{t}/2\right)^{2}\right]},$ (45)
which corresponds to the integrated part in Eq. (8).
Appendix B Each contribution to $B^{\phi\phi\to\phi\phi}_{\text{grav}}$
We show explicit forms of each contribution to $B^{\phi\phi\to\phi\phi}_{\text{grav}}$ in Eq. (24), i.e., $B^{\phi\phi\to\phi\phi}_{\text{grav(a/b/c/d)}}$, corresponding to (a)–(d) in Figure 3, respectively, in this Appendix.
$B^{\phi\phi\to\phi\phi}_{\text{grav(a)}}$ $=-\frac{\lambda^{2}_{h\phi}v^{2}}{96\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f_{(a)}(m_{\phi}/m_{h}),$ (46)
$f_{(a)}(x)$ $=\frac{1}{x^{8}(1-4x^{2})^{2}}\biggl\{120x^{8}-154x^{6}+55x^{4}-6x^{2}-(1-4x^{2})^{2}(3x^{4}-8x^{2}+3)\ln(x^{2})$
$\,\,\,\,\,\,-2\sqrt{1-4x^{2}}(4x^{8}-54x^{6}+69x^{4}-26x^{2}+3)\ln\left[\frac{1}{2x}\left(\sqrt{1-4x^{2}}+1\right)\right]\biggr\},$ (47)
$B^{\phi\phi\to\phi\phi}_{\text{grav(b)}}$ $=-\frac{\lambda^{2}_{h\phi}v^{2}}{96\pi^{2}\overline{M}^{2}_{\text{pl}}m^{4}_{h}}f_{(b)}(m_{\phi}/m_{h}),$ (48)
$f_{(b)}(x)$ $=\frac{1}{x^{8}(1-4x^{2})^{2}}\biggl\{-64x^{8}+124x^{6}-51x^{4}+6x^{2}+(1-4x^{2})^{2}(x^{4}-6x^{2}+3)\ln(x^{2})$
$\,\,\,\,\,\,+2\sqrt{1-4x^{2}}(-30x^{6}+55x^{4}-24x^{2}+3)\ln\left[\frac{1}{2x}\left(\sqrt{1-4x^{2}}+1\right)\right]\biggr\},$ (49)
$B^{\phi\phi\to\phi\phi}_{\text{grav(c)}}$ $=B^{\phi\phi\to\phi\phi}_{\text{grav(a)}},\,\,\,\,B^{\phi\phi\to\phi\phi}_{\text{grav(d)}}=B^{\phi\phi\to\phi\phi}_{\text{grav(b)}}.$ (50)